OpenStudy (anonymous):

trying to figure out 3000 meters in 10 min. use 2 unit multipliers to convert average speed to kilometers per hour

OpenStudy (mertsj):

\[\frac{3000 m}{10 \min}\times \frac{1 km}{1000m}\times \frac{60 \min}{1 hr}\]

OpenStudy (mertsj):

Now if you just cancel the units and do the arithmetic, you should be good.

OpenStudy (anonymous):

Thank you, I will try... Greatly appreciated.

OpenStudy (mertsj):

yw